OpenStudy (anonymous):

A newly discovered solar system is found 300 billion kilometers from the sun. How many AU does this represent? Show your work. PLEASE HELP!

OpenStudy (phoenixfire):

@onaogh That is not going to help people learn. And plus gingy_fizzy is asking Kilometers. @gingy_fizzy If you know how much 1 AU is in kilometers you can simply divide 300 billion km by 1 AU in km and that will give you the amount of AU there is.

OpenStudy (anonymous):

@PhoenixFire i will gladly help those who wants to learn. But those who just want their homework answer, i refer to google. maybe u can do her/his homework

OpenStudy (theeric):

@onaogh , it seems like you have good intentions, but you should watch out for giving answers, because it's again the Code of Conduct. Also, you found 300 billion meters in astronomical units, not 300 billion kilometers. Note that \(1\ [km]=6.68458712 \times 10^{-9}\ [AU]\) according to Google. So, multiply both sides by \(1\ [km]\) to get \(\dfrac{\cancel{1\ [km]}}{\cancel{1\ [km]}}=\dfrac{6.68458712 \times 10^{-9}\ [AU]}{1\ [km]}=1\). And anything multiplied by one is itself! So, if you have \(3.00\times 10^{11}\ [km]\), then \((3.00\times 10^{11}\ [km])(1)=(3.00\times 10^{11}\ [km])\left(\dfrac{6.68458712 \times 10^{-9}\ [AU]}{1\ [km]}\right)\)\(=(3.00\times 10^{11}\ \cancel{[km]})\left(\dfrac{6.68458712 \times 10^{-9}\ [AU]}{1\ \cancel{[km]}}\right)\) As you see, the value will be the same and the units will be astronomical units.

OpenStudy (theeric):

That's a common method for converting units :)